OpenStudy (anonymous):

Compute: \[\log_2 (\frac{2012!!}{1006!})\] NO CALCULATORS PERMITTED!

Parth (parthkohli):

\[{2012 \over 1006} = 2 \]

Parth (parthkohli):

\[ \log_2 2 = ?\]

Parth (parthkohli):

Unless that is really a factorial.

OpenStudy (anonymous):

Yup...By definition: \[n! = n!(n - 1)!(n - 2)!...1!\]\[n!! = n!(n - 2)!(n - 4)!...1! or 2!\]based on whether it's even or odd.

Parth (parthkohli):

That's sad.

OpenStudy (anonymous):

Oops. No ! marks in between lol

OpenStudy (anonymous):

Take 2 common in the numerator.. \(2^{1006}\) (1006)! May be this will help..

OpenStudy (anonymous):

I'm not quite understanding what you mean?

OpenStudy (anonymous):

I show you..

OpenStudy (anonymous):

\[2012!! = (2012)(2010)(2008)(2006)(2004)....................2!\] As these are in multiplication so you have to take 2 common from all the terms: \[2012!! = 2^{1006}(1006)(1005)(1004)(1003)(1002)....................1!\] \[2012!! = 2^{1006} \times 1006!\]

OpenStudy (anonymous):

I am having doubt that should be 1006 or 1005 in the power..

OpenStudy (anonymous):

Oh...so you're dividing each of the parts of the factorial by 2?

OpenStudy (anonymous):

I am taking common (distributive property)..

OpenStudy (anonymous):

So then: \[\log_2(\frac{2012!!}{1006!}) \implies \log_2(\frac{2^{1006} \times 1006!}{1006!}) \implies \log_2(2^{1006}) \implies 1006\]Is that what happens?

OpenStudy (anonymous):

(2012)(2010)(2008) Take 2 common from first: 2(1006) (2010) (2008) Like wise take 2 common from all terms: 2(1006) 2(1005) 2(1004) -------------2^3 (1006)(1005)(1004)

OpenStudy (anonymous):

Like wise..

OpenStudy (anonymous):

It's because all of the terms are even, so you can remove a 2 from each which when multiplied to gether form \(2^{1006}\)since there are 1006 terms in 2012!?

OpenStudy (anonymous):

*factor out

OpenStudy (anonymous):

Yes that is what I am saying.. But I am in a doubt that will be 1005 or 1006..

OpenStudy (anonymous):

Let me think for two minutes..

OpenStudy (anonymous):

I think it would be 1006 since the biggest term 2012, when divided by 2 forms 1006 which the rest follow suit...plus, if you know the context of this problem, there wouldn't be a way to get the end result to be pretty...

OpenStudy (anonymous):

Yes it will be 1006 using AP..

OpenStudy (anonymous):

Wait, what does AP stand for?

OpenStudy (anonymous):

2 + 4 + 6 + 2012 Find n using Arithmetic Progression (A.P).

OpenStudy (anonymous):

They are total 1006 terms so use will have factor out or divide 2 1006 times..

OpenStudy (anonymous):

Ah ok...thank you! I understand it now! :)

OpenStudy (anonymous):

Welcome dear..